OpenStudy (anonymous):

Hey I need some help, I think I know this answer but not completely sure... In your lab, a substance's temperature has been observed to follow the function T(x) = (x − 2)^3 + 8. The turning point of the graph is where the substance changes from a solid to a liquid. Using complete sentences in your written answer, explain to your fellow scientists how to find the turning point of this function. Hint: The turning point of the graph is similar to the vertex of a quadratic function. Since the equation is already in vertex form (I believe) would the vertex just be (2, 8)?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@Nnesha

OpenStudy (baru):

you are right

OpenStudy (anonymous):

@baru when explaining it do I just mention how it is already in vertex form I can infer that it is the vertex. I don't have any like equation to do right?

OpenStudy (baru):

you can graph the function here and double check https://www.desmos.com/calculator

OpenStudy (anonymous):

@baru thank you man!

OpenStudy (baru):

(T-8) = (x − 2)^3 technically that is the vertex form the vertex form is (in general) (y-k)= f(x-h) where (h,k) is the vertex

OpenStudy (anonymous):

ok thank you, I didn't know that was a possible vertex form but I have it written down, as the one I learned was y=a(x-h)^2+k, where (h, k) is the vertex, but thank you again @baru :D

OpenStudy (baru):

both are really the same-thing...follow whichever you are more comfortable with :)